OpenStudy (anonymous):

a building in a city has a rectangular base. the length of the base measures 70 ft less than three times the width. the perimeter of this base is 940 ft. what are the dimensions of the base?

OpenStudy (texaschic101):

Perimeter of rectangle is : P = 2L + 2W P = 940 L = 3W - 70 now we sub 940 = 2(3W - 70) + 2W --- distribute through the parenthesis 940 = 6W - 140 + 2W -- combine like terms 940 = 8W - 140 --- add 140 to both sides 940 + 140 = 8W 1080 = 8W -- divide both sides by 8 1080/8 = W 135 = W L = 3W - 70 L = 3(135) - 70 L = 405 - 70 L = 335 check.. P = 2L + 2W 940 = 2(335) + 2(135) 940 = 670 + 270 940 = 940 (correct) Length (L) = 335 Width (W) = 135